OpenStudy (anonymous):

What is the inverse of the equation y = 5x - 1

Directrix (directrix):

>inverse of the equation y = 5x - 1 Exchange y and x to get: x = 5y - 1 Then, solve for y. Post what you get. Okay?

OpenStudy (anonymous):

y=x+3 5

OpenStudy (anonymous):

y=x+1 5

Directrix (directrix):

x = 5y - 1 5y - 1 = x 5y - 1 + 1 = x + 1 5y = x + 1 Use the reciprocal approach; a number and its reciprocal multiply to 1 (1/5) * (5y) = (1/5)* x + (1/5)* 1 y = (x/5) + (1/5) You might want to leave it as y = (1/5)*(x + 1) --> Where did you get the 3? Just wondering.

OpenStudy (anonymous):

i wasnt thinking i know it was 1 but thanks

Directrix (directrix):

Alrighty.